Operations Report – November 30th

Crew 183 + Operations Report 30NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 6 pm

Solar— SOC 90% (Before generator is run at night)

Diesel –  50 %

Propane – 100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 4

Notes on ATVs: nothing to report

Deimos rover used:  yes

Hours:

Beginning charge: 100

Ending charge: 94

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Burn barrel not complete. Turns out you
can’t just poke holes in it with a pickaxe.  We have a plan for
tomorrow.  Also, contractors will be onsite all day tomorrow.

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: nothing to report

Summary of GreenHab operations: Door was opened this afternoon.
Thinned the babies.  Used 4 gallons of water for watering and prepping
the soil for herbs.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Nothing tonight, thank you.

Operations Report – November 29th

Crew 183 + Operations Report 29NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 5:30 pm

Solar— SOC  91% (Before generator is run at night)

Diesel –  50 %

Propane –  100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:  Propane was delivered today.  A leak was
found in the line of the small tank at the director’s house, and it
was fixed.  We began work on the second rover today but then decided
to go to storage and see it we could find better tools.  We found a
socket set that was the right size!  We will get back on it tomorrow.
We also brought a 55 gallon drum from town.  Tomorrow I will pound
holes in it and set it up as a burn barrel.  I will also need to clean
up the old burn barrel area as non-burnable garbage was burned and it
will need to be picked up and taken for disposal.  Lucky me!

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Three new smoke detectors arrived today.

Summary of GreenHab operations: Door was opened this afternoon.  Two
gallons of water was used in the new planters.  I did not water the
plants today.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Thanks for all of
the help with the rover installation question last night.  No
questions tonight.

Operations Report – November 28th

Crew 183 + Operations Report 28NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 6 pm

Solar— SOC  92% (Before generator is run at night)

Diesel –  50 %

Propane –  not known psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –  not known gallons.

Water (trailer) –  0 gallons.

Water (static) –   300 gallons

Trailer to Static Pump used –  no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  yes

Hours: not recorded (but crews need to record this information)

Beginning charge: 100

Ending charge: 95

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: We began putting the new rovers together
today.  I decided that we would construct the cage, but not modify the
seat at this time.  Instead, we will just remove the back seat rest,
which was what the Mars 160 crew wanted to do originally but I
insisted on the modification.  This will still be done on these but at
a later date because it is much more secure.  Maybe when Yusuke is
here with his crew, since he modified them last time and will be able
to help me to do it.  In the meantime, crews will still be able to use
them.  They will require about a month of short trips before they are
able to go on long ones like Deimos.

Summary of internet:  not known

Summary of suits and radios:  Ashton has recommended that all the old
radios that are sketchy be taken out of service.  I agree.  Four new
radios and headsets will be added to the Hab before the next crew
arrives.  Old sketchy ones will be removed.

Summary of Hab operations: Two new CO alarms were delivered to the
Hab.  We will mount one each on the lower and upper deck before the
new crew arrives.

Summary of GreenHab operations:  Two new planters were added to the
GreenHab.  They are for an herb garden and a greens garden.  I am
prepping the soil.  I turned the heater down to 55 degrees F.  Today I
left the door open all day.  Used 6 gallons of water for plants and
planters.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support:  We have bolts
that have 10.8 and 8.8 on them.  What does that mean in turns of
socket size, if anything?  We don’t have a socket set that fits the
bolts on the rovers, so are trying to figure out if they are metric or
English system.  Any information appreciated.

Operations Report – November 27th

EVA #5 Report

Today our EVA was conducted out-of-sim, with the purpose of further testing our 3D scanning and imagery technology out on the field. Jordan, Ashton, Toruwa, Dean, and I took Deimos and three ATVs to Lith Canyon. We left the hab at approximately 12:20 PM. On the way, we stopped by the location where we obtained the specimens and return them as we found them before.

After spending some time there, we arrived to the Lith Canyon turnout by around 1:15 PM. We walked to the ridge overlooking the canyon and conducted our activities in this area. We stayed here until about 2:30 PM, and made our way back to the Hab by 3:10 PM.

Have a good night Mission Support!

Brandon Ferguson

EVA Officer

Operations Report – November 26th

Crew 183 Operations Report 26-11-2017

SOL: 4

Name of person filing report: Ashton Stoop

Non-nominal systems:

N/A

Notes on non-nominal systems:

Generator (hours run): 14

Solar— 75%

Diesel – 54%

Propane – 40% volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 8 gallons remaining

Water (trailer) – 0 gallons added to static

Water (static) – 300 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – yes

Water Meter: 44

Toilet tank emptied: Yes

ATV’s Used: N/A

Oil Added? No

ATV Fuel Used: 0

# Hours the ATVs were Used today: 0

Notes on ATVs: N/A

Deimos rover used: Yes

Hours: 1

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? NO

General notes and comments:

Summary of internet:

Every evening after dinner time it slows down a lot.

Summary of suits and radios:

A lot of the radios have been giving us issues. Sometimes they work and sometimes they don’t. It seems to be a bit random which is a problem when we are trusting them out on 3-hour EVAs. A few of the suits aren’t charging properly. One of the suits failed today during EVA and we had to rush back to the hab before a crew member’s suit was compromised.

Summary of Hab operations:

Things are operating smoothly.

Summary of GreenHab operations:

Everything is growing well.

Summary of ScienceDome operations:

Sedementation is progressing well.

Summary of RAMM operations:

Not used. It is not yet fully operational.

Summary of health and safety issues:

Waiting and Carbon monoxide detectors.

Questions, concerns and requests to Mission Support:

-There definitely needs to be some further diagnostics performed on suits and radios.

Operations Report – November 25th

Crew 183 Operations Report 25-11-2017

SOL: 3

Name of person filing report: Ashton Stoop

Non-nominal systems:

N/A

Notes on non-nominal systems:

Generator (hours run): 12

Solar— 85%

Diesel – 60%

Propane – 41% volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 8 gallons remaining

Water (trailer) – 0 gallons added to static

Water (static) – 325 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – yes

Water Meter: 48

Toilet tank emptied: No

ATV’s Used: Honda, 350.1

Oil Added? No

ATV Fuel Used: 2 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: Everything ran smoothly today.

Deimos rover used: Yes

Hours: 1

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? NO

General notes and comments:

It’s hot in those EVA suits…

Summary of internet:

Every evening after dinner time it slows down a lot.

Summary of suits and radios:

There have been some issues with radios and headsets over the past few days. We will need to do a test and see which ones are working properly and which aren’t. Some of the EVA suits are having issues. Suit 4 only had one fan working and suit 1 wasn’t working at all. We will double check all the charges to see if that was an issue. One of the XLR cables isn’t working properly. We’re going to measure the current going through and give a more detailed diagnosis in the near future.

Summary of Hab operations:

Things are operating smoothly.

Summary of GreenHab operations:

Everything is growing well and grow lights are operational.

Summary of ScienceDome operations:

Sedementation tests are proceeding as planned. We received a new soil solidity device today.

Summary of RAMM operations:

Not used. It is not yet fully operational.

Summary of health and safety issues:

Waiting and Carbon monoxide detectors.

Questions, concerns and requests to Mission Support:

-Things are going well overall.

Operations Report – November 24th

Crew 183 Operations Report 24-11-2017

SOL: 2

Name of person filing report: Ashton Stoop

Non-nominal systems:

Power system – SOC

Notes on non-nominal systems:

The power system calculates SOC based on an algorithm, it is not a measured value. This has been causing strange fluctuations on when reading the SOC.

READING FROM 24 NOV 2017

TIME     SOC READINGS

6:32       61%

9:25       69%

10:11     100%

11:40     100%

16:10     88%

18:37     100%

Generator (hours run): 14

Solar— 100%

Diesel –  62%

Propane –  42% volume

Ethanol Free Gasoline (5 Gallon containers for ATV) –   0 gallons remaining

Water (trailer) – 100 gallons added to static

Water (static) –   325 gallons

Trailer to Static Pump used –  yes

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  45

Toilet tank emptied: yes

ATV’s Used: Honda, 300, 350.1, 350.2

Oil Added? No

ATV Fuel Used:  5 Gals

# Hours the ATVs were Used today: 3

Notes on ATVs: ATV 350.1 had a bit of trouble starting up upon return from EVA. Once it got going everything seemed to be fine. We’ll keep an eye on this for the next few SOLs.

Deimos rover used:   Yes

Hours: 1

Beginning charge: 100%

Ending charge: 90%

Currently charging: Yes

 

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

 

Spirit rover used:  NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

 

Opportunity rover used:  NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? NO

General notes and comments:

The SOC readings seem off. Looking to further investigate this matter.

Summary of internet:

We hit our internet limit around 20:00.

Summary of suits and radios:

Not used

Summary of Hab operations:

Everything is in good shape. We had some bomb crab biscuits for dinner.

Summary of GreenHab operations:

A new lighting timer has been delivered and installed in the GreenHab.

Summary of ScienceDome operations:

New lights have been delivered. Tomorrow, we need to install lightbulbs and test.

Summary of RAMM operations:

Not used. It is not yet fully operational.

Summary of health and safety issues:

Four carbon monoxide detectors have been ordered.

Questions, concerns and requests to Mission Support:

-Dear mission support, how exactly does the SOC reading on the power system operate? We have had fluctuations in the SOC readings which do not correlate closely with our estimated values of power generation and power consumption (Please refer to SOC Readings table above).

-We have no additional supply of gasoline for the ATVs. All that’s left is what’s in the vehicles. We need fuel ASAP!

Operations Report – November 23rd

Operations Report

Crew 183 Operations Report – 23 Nov 2017

SOL: 1

Name of person filing report:  Ashton Stoop

Non-nominal systems:

WATER HEATER: Water heater is still not installed. Hopefully this will be put into place soon.

Generator—  The generator has been turned on at 18:10.

READING FROM 23 NOV 2017

TIME

SOC READINGS

9:20

100%

12:20

100%

15:20

78%

18:10

62%

19:20

62%

20:05

62%

Diesel – 63 %

Oil Pressure: 47.5 PSI

Propane – 43% volume

Gasoline (5 Gallon containers for ATV) – 1 gallons.

Water (trailer) –  0 gallons

Water (static) – 40 gallons

Trailer to Static Pump used –   no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  47 gallons

Toilet tank emptied:  no

ATV’s Used: Yes

Oil Checked:  No

Atv Fuel Used Gals:  3  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 2 hours

Deimos used:  Yes

Hours: 0.5 hours

Beginning charge:  100%

Ending charge:  –

Currently charging: No

HabCar used? No

Notes and Comments:
The door to the Science Dome was not closing properly. The metal latch had been bent and was no longer sliding and latching the door. I was able to bend the metal back into place and the latch is now operational again.

We are having some issues with the power system. At 18:10 the SOC was at 62%. The generator was then turned on, but after 2.6 hours of using the generator, the SOC is still at 62%. My first thought is simply that it’s not reading correctly, but I am slightly concerned that it’s not charging properly.

Summary of GreenHab operations:

The GreenHab is looking in pretty good shape. There is a red wheelbarrow which has yet to be assembled. The axel is long enough to house the wheels, and all the parts are present, however there are protruding metal sleeves covering part of the axel. If this metal could be removed, then the wheels would easily slide on.

Summary of ScienceDome operations:

N/A

Summary of health and safety issues:

The carbon monoxide detector is not working, even with new batteries it won’t turn on.

Questions, concerns and requests to Mission Support:

-We are going to need gasoline for the ATVs ASAP.

-We need to figure out this whole power system.

Operations Report – November 22nd

Crew 182+ Operations Report  22NOV2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Lower deck wall heater is now operational, but
forced air furnace continues to alternate between warm and cool air.

I keep forgetting to report than the porthole came out of the main
airlock again.

Notes on non-nominal systems: nothing to report

Generator (hours run): Ran until 9 am, then off until 5:45 pm

Solar— SOC 97 % when generator was turned on

Diesel –  not checked %

Propane –  not checked psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –   0 gallons.

Water (trailer) –  0 gallons.

Water (static) –  350 gallons

Trailer to Static Pump used –  no

Water (loft) – Static to Loft Pump used –  not known

Water Meter:  not checked

Toilet tank emptied: Yesterday

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) 350 s moved to their normal place

Oil Added? no

ATV Fuel Used:  not known Gals

# Hours the ATVs were Used today: 5 min

Notes on ATVs: The 300’s battery is low, did not jump today.

Deimos rover used:  yes

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR, not yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used:  Not yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Opportunity rover used:  not yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used:  no yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: no

HabCar used? Yes, hauled 100 gallons of water and emptied the 100
gallons Brandon got yesterday

General notes and comments: Cleaning day

Summary of internet: not known

Summary of suits:  I checked all of them today.  Two more chargers not
functional.  Unfortunately one of the helmets was cracked in such a
way that it will not be repairable.  Removed from service.  One of the
backpacks is missing one of the two straps that hold the collar on the
frame.

Summary of Hab operations: Brandon and I cleaned the Hab.  It took all
morning and then some.  Looking good now.

Summary of GreenHab operations:  Watered.  Used 4 gallons.  Removed
hose from frost free faucet.  Crews should get water from faucet
without hose from now until spring to prevent freeze up.  I’ve been
thinking I may get the 55 gallon drum from town to keep water in
inside the GreenHab.  It might also increase humidity, which seems to
stay around 16 %.

Summary of ScienceDome operations: Nothing to report.

Summary of RAMM operations:  Nothing to report.

Summary of health and safety issues:  We are safe and we are healthy.

Questions, concerns and requests to Mission Support:  Nothing, thank you.

Operations Report – November 20th

Crew 182+ Operations Report  20NOV2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: same as yesterday, except the generator is
running well again

Notes on non-nominal systems:  Generator run all night last night, so
we don’t know if the last failure was due to the diesel or needing
oil.

Generator (hours run): 24

Solar— SOC 100 %

Diesel –  63 %

Propane –  not known psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –   0 gallons.

Water (trailer) –  0 gallons.

Water (static) –  150 gallons (Brandon is currently bringing 100
gallons but we will need to make a couple more trips tomorrow.

Trailer to Static Pump used –  no

Water (loft) – Static to Loft Pump used –  not known

Water Meter:  not checked

Toilet tank emptied: not that I know of

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) All 5 were moved to make
space for the big equipment preparing for the RAMM’s arrival tomorrow

Oil Added? no

ATV Fuel Used:  0Gals

# Hours the ATVs were Used today: 5 min each

Notes on ATVs: The 300’s battery is low, will need to jump it tomorrow

Deimos rover used:  yes

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR, not yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used:  Not yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Opportunity rover used:  not yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used:  no yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: no

HabCar used? Yes

General notes and comments:  The new rovers need to be put together
and the seats modified (except for Sojourner).  Not yet started on
this project.  Will charge the other three tomorrow.

Summary of internet:  Internet was upgraded at Shannon’s, but the
phone does not work properly.  Internet seems as good as before it
went bad, but not blindingly faster even with the upgrade.  Wifi is
turned off to this service.  Hab internet is good.

Summary of suits:  n/a, not used

Summary of Hab operations: The forced air furnace is still
periodically blowing cold air.  Need to clean the filter tomorrow then
order some as soon as possible to see if that is the problem.

Summary of GreenHab operations:  I moved the beans to the wooden bench
and provided cages for their continued growth.  Other plants were
organized.  I prepped 10 pots for moringa seeds.  Ordered further
supplies: timer for lights, environmental monitor, herb and lettuce
seeds with planters

Summary of ScienceDome operations: Ordered two floor lamps for evening
light.  We will have to clean the lab tomorrow as it is not in
changeover condition.

Summary of RAMM operations:  The area was leveled and the swale was
dug so that water will go around the building during rain and runoff.
The RAMM left California this evening and will be onsite at 10 am.  It
will be lowered onto a railroad tie frame to keep it off the ground.

Summary of health and safety issues:  Nothing to report

Questions, concerns and requests to Mission Support:  Nothing, thank you.